OpenStudy (anonymous):

Sarah is shopping for carpet in her square dorm room. The carpet advertised is 121 square feet. Which expression represents the length of one side of Sarah’s square room?

OpenStudy (stamp):



OpenStudy (stamp):

\[A_{square}=s^2\]\[A=121ft^2\]\[121ft^2=s^2\]\[s=11ft\]